# End-to-End Protocol for High-Quality QAOA Parameters with Few Shots

### v1

@@ Abstract
The quantum approximate optimization algorithm (QAOA) is a quantum heuristic for combinatorial optimization that has been demonstrated to scale better than state-of-the-art classical solvers for some problems. For a given problem instance, QAOA performance depends crucially on the choice of the parameters. While average-case optimal parameters are available in many cases, meaningful performance gains can be obtained by fine-tuning these parameters for a given instance. This task is especially challenging, however, when the number of circuit executions (shots) is limited. In this work, we develop an end-to-end protocol that combines multiple parameter settings and fine-tuning techniques. We use large-scale numerical experiments to optimize the protocol for the shot-limited setting and observe that optimizers with the simplest internal model (linear) perform best. We implement the optimized pipeline on a trapped-ion processor using up to $32$ qubits and $5$ QAOA layers, and we demonstrate that the pipeline is robust to small amounts of hardware noise. To the best of our knowledge, these are the largest demonstrations of QAOA parameter tuning on a trapped-ion processor.
@@ I Introduction
The optimization of QAOA parameters presents a significant challenge if the number of shots is limited. This challenge is exacerbated by the fundamental limits that quantum mechanics imposes on the cost of computing gradients of quantum circuits [25]. The high cost of computing the gradient motivates the use of derivative-free optimization, which typically either assumes a deterministic objective [26] or requires a high number of shots to converge [27, 28, 29, 30]. As a consequence of these challenges and despite the recent progress [31, 32, 33, 34, 35, 36, 37, 38, 39, 35], the problem of optimizing parameterized quantum circuits with a small number of shots remains open.
@@ III.2 Optimizer choice
We conduct an evaluation of various optimization methods under the shot-frugal setting. Specifically, we examine the performance of COBYLA [42], BOBYQA [43], NELDER_MEAD [44], ESCH [45], DIRECT_L [46], CRS2_LM [47], SPSA [48], GSLS [49], and IMFIL [50]. These methods have been considered in the quantum optimization context and are available in optimization packages NLopt [51], SciPy [52], PDFO [53], and Scikit-Quant [54]. Some other methods we also tested but either work very similarly to one of the above methods or perform undesirably include UOBYQA [55], NEWUOA [56], LINCOA [57], and SNOBFIT [58]. We compute the energy landscapes of 60 random $p$=1 PO instances and efficiently test the optimization quality of each method.
@@
Figure 2 shows the performance comparison of the tested methods using mean relative AR improvement (eq. 12) as the metric. We see that BOBYQA, COBYLA, and NELDER_MEAD perform the best in this setting among all tested methods. We choose COBYLA over NELDER_MEAD since the former is considered as an improved version of the latter [42]. Between BOBYQA and COBYLA, we choose the latter, attributing to the fact that a simple model minimizes the number of initial function evaluations and maximizes the number of shots per evaluation. COBYLA assumes a linear model and needs only $12p+$ 1 initial function evaluations to build the model for $2p$ parameters in a $p$-layer QAOA. BOBYQA follows almost the same strategy as COBYLA except that it assumes a quadratic model, which requires $\frac{1}{2}(2p+1)(2p+2)$ initial function evaluations to fully determine the model. This quadratic growth is less desirable than COBYLA’s linear growth as we go to higher $p$s. Similarly, other methods generally have more complicated assumptions and are thus more demanding in terms of the number of evaluations. The improved accuracy in the predictions does not compensate for the loss resulting from the large decrease in the number of shots per evaluation.
@@ III.3 Hyperparameter selection
We now fix COBYLA as our optimization method and investigate its hyperparameter choice. COBYLA only has one hyperparameter to tune besides the budget allocation strategy, which is the initial step size (“rhobeg”). As previously mentioned, this hyperparameter determines the distance between the initial point and other initial evaluations, which is crucial for establishing an accurate linear model while using few evaluations. We run exact simulations of 100 random 12-qubit MaxCut instances and 60 random 12-qubit PO instances from $p=1$ to 5 and plot the optimization results under varying initial step sizes.
In Figure 3, we plot the mean relative AR improvement with standard error over instances as a function of initial step size, where each curve represents the performance at the given number of function evaluations after the initial evaluations. The performance converges toward the maximum possible relative improvement as we increase the function evaluations, as expected. We observe that MaxCut, having a more accurate initialization strategy, requires a smaller initial step size than does PO. With a given problem and initialization strategy, we notice that COBYLA is generally not sensitive to the initial step size, even when the number of evaluations is very low. The same trend can be observed across different $p$s, which we show in Figure 11 and Figure 12 in Appendix A. For subsequent experiments, we use an initial step size of 0.1 for MaxCut and 0.5 for PO.
@@ III.4 Budget allocation
We now study the budget allocation strategy. We primarily focus on a per-instance total budget of $10,000$ shots. This value of total budget was chosen to match the constraints of trapped-ion hardware. Prior QAOA experiments on the device used in this work used $<5000$ shots and $p\leq 2$ [22]. For comparison, previous hardware demonstration on a superconducting processor used 25,000 shots per evaluation and 6 evaluations per optimizer iteration [59]. The number of shots per evaluation is inversely proportional to the number of total function evaluations:
@@
We perform hyperparameter grid searches on the number of shots per function evaluation versus the maximum number of function evaluations on 1,000 random 12-qubit MaxCut instances from $p=2$ to 5.
@@
Figure 5 shows a contour plot spanned by the number of additional evaluations and the number of shots per evaluation. The color represents mean relative AR improvement, and the three lines correspond to a total budget of 10k, 20k, and 30k, respectively. Focusing on the 10k budget line, we make the same observation as in Figure 4: smaller number of steps and higher per-evaluation shot budget give the best results. Following the contours, We see that increasing the number of evaluations has little return, in contrast to the steady gain in increasing the number of shots per evaluation. If the number of shots is as low as 200, the optimizer will even find lower-quality parameters than the initial ones due to the extreme variance in sampling the objective function values. We expect the budget allocation strategy to be the same for similar budgets, whereas a substantially higher budget can allow more evaluations. We show results for $p\in\{2,3,4\}$ in Figure 13 and Figure 14 in Appendix A.
@@ III.5 Hardware demonstrations
Figure 6: Optimizing QAOA parameters with $p=1,\ldots,5$ for $n=12$ weighted graphs. Error bars represent the standard error of the mean AR over 5 instances. The Noiseless ()($\infty$)( ∞ ) line optimizes the QAOA parameter under the noiseless backend and with an unlimited number of iterations. In other words, it represents the performance of the best achievable parameters. The Noiseless $(2)$ line optimizes the QAOA parameter under the noiseless backend and with two additional iterations, representing the results of the protocol in noiseless simulation. The H1-1E $(2)$ line optimizes with two additional iterations under an emulator backend , which mimics the H1-1 quantum device, representing the deployment of the protocol in a practical scenario. The Initial line presents the performance under the fixed initial parameter.
Figure 7: Optimizing QAOA parameters with $p=1,3,5$ for three $n=20$ MaxCut instances. Error bars represent the standard error of the mean AR over limited shots. An error bar is estimated as $\frac{\sigma}{\sqrt{M}}$, where $\sigma$ is the standard deviation of the approximation ratio associated with the exact QAOA state vector and $M$ is the number of shots. The added H1-1 (2) line optimizes for 2 additional iterations after the first $12p+$ 1 initial evaluations. For the description of other labels, please refer to the caption of Figure 6.
@@
Figure 9: Optimizing QAOA parameters with $p=1,\ldots,7$ for $n=10$ PO instances. Error bars represent the standard error of the mean AR over $5$ instances. For a description of labels, please refer to the caption of Figure 6.
@@
Table 1: Numerical simulation and hardware demonstration results of $p=5$ QAOA on an $n=32$ MaxCut instance. Values in parentheses are the standard error of the mean AR over limited shots. “2 iter” denotes the results including 2 iterations after the first $12p+$ 1 initial evaluations. The “Hardware evaluation” column shows the results directly reported by the hardware optimization. The “Exact evaluation” column shows the results evaluated with a noiseless state vector simulator using the parameters found by the hardware optimization.
@@ IV Discussion
We propose an end-to-end protocol for QAOA parameter setting. Focusing on weighted combinatorial optimization problems within the same family, we assume we have a fixed parameter initialization schedule for the unweighted problems from the same family. We first rescale the weighted problem instance such that the parameters for unweighted problems can be applicable. Then, we need to tune the hyperparameters of the classical optimizer. Under the shot-limited cases, we benchmarked 12 optimization algorithms and found that COBYLA and BOBYQA perform the best. Considering a realistic shot budget of 10k per problem instance, we use the simplest method, COBYLA. To tune the hyperparameters of COBYLA, we observe that a fixed initial step size works stably in QAOA with different layers for problems of different sizes. Then, we determine the maximum number of iterations of COBYLA to optimize the budget allocation. Under the setting of 10k shots, we assume an equally allocated budget and performe a grid search on the number of shots for one sample. We set the number of steps to be 2 and observe that it consistently works well.
@@
This protocol provides an end-to-end solution for QAOA parameter settings on a quantum device. The protocol is also generally applicable when we have many more shots, where we could run more fine-tuning iterations. For that scenario, some open and interesting directions still exist. For example, we can allocate the shots dynamically, allowing each iteration to have different shots. We can also assign the shot budget to the problem instances that perform relatively poorer under the fixed initial parameters.
@@ V.2 Low-cost optimizer benchmarking
To simulate the energy with finite sampling errors, in addition to the energy landscapes, we compute the standard deviation landscapes to realize fast energy evaluation with an arbitrary shot. Based on the central limit theorem [64], the deviation of a shotted estimation and exact QAOA energy will converge to a Gaussian distribution $\mathcal{N}(0,\frac{\sigma^{2}_{\text{exact}}}{\#\text{shots}})$, where $\sigma_{\text{exact}}$ is the standard derivation of the energy evaluated by an exact QAOA state.
@@ Appendix A Additional results
Here, we show additional results we have obtained, including the initial step size study for $p=1,\ldots,4$ (Figure 11 and Figure 12) and the budget allocation study for $p\in\{2,3,4\}$ (Figure 13 and Figure 14).
(a) $p=1$
(b) $p=2$
(c) $p=3$
(d) $p=4$
Figure 11: Mean relative AR improvement (with standard error over instances) of COBYLA on $p=1,\ldots,4$ MxCut instances as a function of initial step size. The label of each line represents the number of function evaluations allowed after the initial evaluations. We observe that with a given problem and initialization strategy, COBYLA is generally not sensitive to the initial step size or QAOA depth.
(a) $p=1$
(b) $p=2$
(c) $p=3$
(d) $p=4$
Figure 12: Mean relative AR improvement (with standard error over instances) of COBYLA on $p=1,\ldots,4$ PO instances as a function of initial step size. The label of each line represents the number of function evaluations allowed after the initial evaluations. We observe that with a given problem and initialization strategy, COBYLA is generally not sensitive to the initial step size or QAOA depth.
(a) $p=2$
(b) $p=3$
(c) $p=4$
Figure 13: Mean relative AR improvement (with standard error over instances) of optimizing 1,000 $p=2,3,4\ n=12$ MaxCut instances as a function of the number of additional evaluations after the first $12p+$ 1 initial evaluations. Note that for $p=2$, the initial points are of very high quality, and the maximum achievable AR is relatively low due to the low QAOA depth. Consequently, the improvable AR is very small, and the optimizer struggles to optimize beyond the quality of the initial point with a highly stochastic objective resulting from the low shot budget. This is also the reason we do not show $p=1$ figures or use $p=1$ MaxCut in the optimizer benchmarking experiments.
(a) $p=2$
(b) $p=3$
(c) $p=4$
Figure 14: Contour plot of optimizing 1,000 $p=2,3,4\ n=12$ MaxCut instances spanned by the number of additional evaluations and the number of shots per evaluation. The color represents mean relative AR improvement, and the three lines correspond to a total budget of 10k, 20k, and 30k, respectively.

### v3

@@ Abstract
The quantum approximate optimization algorithm (QAOA) is a quantum heuristic for combinatorial optimization that has been demonstrated to scale better than state-of-the-art classical solvers for some problems. For a given problem instance, QAOA performance depends crucially on the choice of the parameters. While average-case optimal parameters are available in many cases, meaningful performance gains can be obtained by fine-tuning these parameters for a given instance. This task is especially challenging, however, when the number of circuit executions (shots) is limited. In this work, we develop an end-to-end protocol that combines multiple parameter settings and fine-tuning techniques. We use large-scale numerical experiments to optimize the protocol for the shot-limited setting and observe that optimizers with the simplest internal model (linear) perform best. We implement the optimized pipeline on a trapped-ion processor using up to $32$ qubits and $5$ QAOA layers, and we demonstrate that the pipeline is robust to small amounts of hardware noise. To the best of our knowledge, these are the largest demonstrations of QAOA parameter fine-tuning on a trapped-ion processor in terms of 2-qubit gate count.
@@ I Introduction
The optimization of QAOA parameters presents a significant challenge if the number of shots is limited. This challenge is exacerbated by the fundamental limits that quantum mechanics imposes on the cost of computing gradients of quantum circuits [25]. The high cost of computing the gradient motivates the use of derivative-free optimization, which typically either assumes a deterministic objective [26] or requires a high number of shots to converge [27, 28, 29, 30]. As a consequence of these challenges and despite the recent progress [31, 32, 33, 34, 35, 36, 37, 38, 39, 35, 40, 41], the problem of optimizing parameterized quantum circuits with a small number of shots remains open.
@@ III.2 Optimizer choice
We conduct an evaluation of various optimization methods under the shot-frugal setting. Specifically, we examine the performance of COBYLA [44], BOBYQA [45], NELDER_MEAD [46], ESCH [47], DIRECT_L [48], CRS2_LM [49], SPSA [50], GSLS [51], and IMFIL [52]. These methods have been considered in the quantum optimization context and are available in optimization packages NLopt [53], PyBOBYQA [54], SciPy [55], PDFO [56], and Scikit-Quant [57]. Some other methods we also tested but either work very similarly to one of the above methods or perform undesirably include UOBYQA [58], NEWUOA [59], LINCOA [60], and SNOBFIT [61]. We compute the energy landscapes of 60 random $p$=1 PO instances and efficiently test the optimization quality of each method.
@@
Figure 2 shows the performance comparison of the tested methods using mean relative AR improvement (eq. 12) as the metric. We see that BOBYQA, COBYLA, and NELDER_MEAD perform the best in this setting among all tested methods. We choose COBYLA over NELDER_MEAD since the former is considered an improved version of the latter [44]. Between BOBYQA and COBYLA, we choose the latter, attributing to the fact that a simple model minimizes the number of initial function evaluations and maximizes the number of shots per evaluation. COBYLA assumes a linear model and needs only $12p+$ 1 initial function evaluations to build the model for $2p$ parameters in a $p$-layer QAOA. BOBYQA follows almost the same strategy as COBYLA except that it assumes a quadratic model, which requires up to $\frac{1}{2}(2p+1)(2p+2)$ initial function evaluations to fully determine the model. We adopt the default setting of $14p+$ 1 initial evaluations, which is still almost twice as many as COBYLA requires in the large $p$ limit. In the small $p$ regions, the performance of BOBYQA and COBYLA are comparable (see more details in Appendix A). However, the number of shots per evaluation for BOBYQA approaches half of that for COBYLA as we go to higher $p$s. The substantial uncertainty due to shotted evaluations outweighs the benefits of fitting a more refined model. Similarly, other methods generally have more complicated assumptions and are thus more demanding in terms of the number of evaluations. The improved accuracy in the predictions does not compensate for the loss resulting from the significant decrease in the number of shots per evaluation.
@@ III.3 Hyperparameter selection
We now fix COBYLA as our optimization method and investigate its hyperparameter choice. COBYLA only has one hyperparameter to tune besides the budget allocation strategy, which is the initial step size (“rhobeg”). As previously mentioned, this hyperparameter determines the distance between the initial point and other initial evaluations, which is crucial for establishing an accurate linear model while using few evaluations. We run exact simulations of 100 random 12-qubit MaxCut instances and 60 random 12-qubit PO instances for $p\in\{1,\ldots,5\}$ and plot the optimization results under varying initial step sizes.
In Figure 3, we plot the mean relative AR improvement with standard error over instances as a function of initial step size, where each curve represents the performance at the given number of function evaluations after the initial evaluations. The performance converges toward the maximum possible relative improvement as we increase the function evaluations, as expected. We observe that MaxCut, having a more accurate initialization strategy, requires a smaller initial step size than does PO. With a given problem and initialization strategy, we notice that COBYLA is generally not sensitive to the initial step size, even when the number of evaluations is very low. The same trend can be observed across different $p$s, which we show in Figure 11 in Appendix A. For subsequent experiments, we use an initial step size of 0.1 for MaxCut and 0.5 for PO.
@@ III.4 Budget allocation
We now study the budget allocation strategy. We primarily focus on a per-instance total budget of $10,000$ shots. This value of total budget was chosen to match the constraints of trapped-ion hardware. Prior QAOA experiments on the device used in this work used 2000, 5200, and 7800 shots and $p\leq 2$ [62, 22]. For comparison, previous hardware demonstration on a superconducting processor used 25,000 shots per evaluation and 6 evaluations per optimizer iteration [63]. The number of shots per evaluation is inversely proportional to the number of total function evaluations:
@@
We perform hyperparameter grid searches on the number of shots per function evaluation versus the maximum number of function evaluations on 1,000 random 12-qubit MaxCut instances for $p\in\{2,3,4,5\}$.
@@
Figure 5 shows a contour plot spanned by the number of additional evaluations and the number of shots per evaluation. The color represents mean relative AR improvement, and the three lines correspond to a total budget of 10k, 20k, and 30k, respectively. Focusing on the 10k budget line, we make the same observation as in Figure 4: smaller number of steps and higher per-evaluation shot budget give the best results. Following the contours, We see that increasing the number of evaluations has little return, in contrast to the steady gain in increasing the number of shots per evaluation. If the number of shots is as low as 200, the optimizer will even find lower-quality parameters than the initial ones due to the extreme variance in sampling the objective function values. We expect the budget allocation strategy to be the same for similar budgets, whereas a substantially higher budget can allow more evaluations. We show additional results for $p\in\{2,3,4\}$ in Figure 12 and Figure 13 in Appendix A.
@@ III.5 Hardware demonstrations
Figure 6: Optimizing QAOA parameters with $p\in\{1,\ldots,5\}$ for $n=12$ weighted graphs. Error bars represent the standard error of the mean AR over 5 instances. The Noiseless ()($\infty$)( ∞ ) line optimizes the QAOA parameter under the noiseless backend and with an unlimited number of iterations. In other words, it represents the performance of the best achievable parameters. The Noiseless $(2)$ line optimizes the QAOA parameter under the noiseless backend and with two additional iterations, representing the results of the protocol in noiseless simulation. The H1-1E $(2)$ line optimizes with two additional iterations under an emulator backend , which mimics the H1-1 quantum device, representing the deployment of the protocol in a practical scenario. The Initial line presents the performance under the fixed initial parameter.
Figure 7: Optimizing QAOA parameters with $p\in\{1,3,5\}$ for three $n=20$ MaxCut instances. Error bars represent the standard error of the mean AR over limited shots. An error bar is estimated as $\frac{\sigma}{\sqrt{M}}$, where $\sigma$ is the standard deviation of the approximation ratio associated with the exact QAOA state vector and $M$ is the number of shots. The added H1-1 (2) line optimizes for 2 additional iterations after the first $12p+$ 1 initial evaluations. For the description of other labels, please refer to the caption of Figure 6.
@@
Figure 9: Optimizing QAOA parameters with $p\in\{1,\ldots,7\}$ for $n=10$ PO instances. Error bars represent the standard error of the mean AR over $5$ instances. For a description of labels, please refer to the caption of Figure 6.
@@
Table 1: Numerical simulation and hardware demonstration results of $p=5$ QAOA on an $n=32$ MaxCut instance. Values in parentheses are the standard error of the mean AR over limited shots. “2 iter” denotes the results, including 2 iterations after the first $12p+$ 1 initial evaluations. The “Hardware evaluation” column shows the results directly reported by the hardware optimization. The “Exact evaluation” column shows the results evaluated with a noiseless state vector simulator using the parameters found by the hardware optimization.
@@ IV Discussion
We propose an end-to-end protocol for QAOA parameter setting. Focusing on weighted combinatorial optimization problems within the same family, we assume we have a fixed parameter initialization schedule for the unweighted problems from the same family. We first rescale the weighted problem instance such that the parameters for unweighted problems can be applicable. Then, we need to tune the hyperparameters of the classical optimizer. Under the shot-limited cases, we benchmark 12 optimization algorithms and found that COBYLA and BOBYQA perform the best. Considering a realistic shot budget of 10k per problem instance, we use the simplest method, COBYLA. To tune the hyperparameters of COBYLA, we observe that a fixed initial step size works stably in QAOA with different layers for problems of different sizes. Then, we determine the maximum number of iterations of COBYLA to optimize the budget allocation. Under the setting of 10k shots, we assume an equally allocated budget and perform a grid search on the number of shots for one sample. We set the number of steps to be 2 and observe that it consistently works well.
@@
This protocol provides an end-to-end solution for QAOA parameter settings on a quantum device. The protocol is also generally applicable when we have many more shots, allowing for more fine-tuning iterations. In such scenarios, several open and interesting directions remain. For example, we can dynamically allocate the shots, allowing each iteration to have a different number of shots. Additionally, we can assign the shot budget to problem instances that perform relatively poorly under the fixed initial parameters. Finally, we remark that our techniques can be combined with other qubit-reuse compilation techniques (e.g., Ref. [62]) to dramatically increase the size of the problem that can be tackled.
@@ V.2 Low-cost optimizer benchmarking
To simulate the energy with finite sampling errors, in addition to the energy landscapes, we compute the standard deviation landscapes to realize fast energy evaluation with an arbitrary shot. Based on the central limit theorem [68], the shotted estimation of the QAOA energy converges to a Gaussian distribution $\mathcal{N}(E,\frac{\sigma^{2}}{\#\text{shots}})$, where $E$ and $\sigma$ are the mean and standard deviation of the energy evaluated by an exact QAOA state.
@@ Appendix A Additional results
Here, we show additional results we have obtained, including the initial step size study for $p\in\{1,2,3,4\}$ (Figure 11) and the budget allocation study for $p\in\{2,3,4,5\}$ (Figure 12 and Figure 13).
(a) $p=1$ MaxCut
(b) $p=2$ MaxCut
(c) $p=3$ MaxCut
(d) $p=4$ MaxCut
(e) $p=1$ PO
(f) $p=2$ PO
(g) $p=3$ PO
(h) $p=4$ PO
Figure 11: Mean relative AR improvement (with standard error over instances) of COBYLA on $p\in\{1,2,3,4\}$ MaxCut and PO instances as a function of initial step size. The label of each line represents the number of function evaluations allowed after the initial evaluations. We observe that with a given problem and initialization strategy, COBYLA is generally not sensitive to the initial step size or QAOA depth.
In Figure 12, following the similar setup as Figure 4, we did an additional comparison between COBYLA implemented in NLopt [53] and Py-BOBYQA [54]. The contours of their performance under different budgets are shown in Figure 13. Py-BOBYQA is an improved version of BOBYQA implemented in Python. In our numerical experiments, we found that the performance of Py-BOBYQA is similar to the standard BOBYQA implemented in NLopt [53]. We did not enable the objfun_has_noise flag in Py-BOBYQA due to its effect of defaulting to $\frac{1}{2}(2p+1)(2p+2)$ initial function evaluations (instead of $14p+$ 1) and using multi-restarts, both of which contradicted our shot-frugal setting. We also wanted to be consistent with the $14p+$ 1 initial function evaluations we used for BOBYQA in our optimizer comparison experiment (Figure 2).
Note that for $p=1$, the initial points are of very high quality, and the maximum achievable approximation ratio (AR) is relatively low due to the shallow QAOA depth. Consequently, the improvable AR is very small, and the optimizer struggles to improve beyond the quality of the initial point, especially with a highly stochastic objective resulting from the low shot budget. Therefore, we do not show $p=1$ figures or use $p=1$ MaxCut in the optimizer benchmarking experiments.
For a small $p$, the difference in the number of initial evaluations does not result in a huge disparity in the number of shots per evaluation. At $p=2$, Py-BOBYQA outperforms COBYLA with its quadratic model. For $p\in\{3,4,5\}$, their performances are comparable, and COBYLA shows a progressive momentum.
(a) COBYLA $p=2$
(b) COBYLA $p=3$
(c) COBYLA $p=4$
(d) COBYLA $p=5$
(e) Py-BOBYQA $p=2$
(f) Py-BOBYQA $p=3$
(g) Py-BOBYQA $p=4$
(h) Py-BOBYQA $p=5$
Figure 12: Mean relative AR improvement (with standard error over instances) of optimizing 1,000 $p\in\{2,3,4,5\}\ n=12$ MaxCut instances as a function of the number of additional evaluations after the first $12p+$ $14p+$ 1 initial evaluations for COBYLA/Py-BOBYQA.
(a) COBYLA $p=2$
(b) COBYLA $p=3$
(c) COBYLA $p=4$
(d) COBYLA $p=5$
(e) Py-BOBYQA $p=2$
(f) Py-BOBYQA $p=3$
(g) Py-BOBYQA $p=4$
(h) Py-BOBYQA $p=5$
Figure 13: Contour plot of optimizing 1,000 $p\in\{2,3,4,5\}\ n=12$ MaxCut instances spanned by the number of additional evaluations and the number of shots per evaluation. The color represents mean relative AR improvement, and the three lines correspond to a total budget of 10k, 20k, and 30k, respectively.